### v1

@@ 3.3.2 Architectural Difference
Traditional models offer a straightforward approach to toxicity detection in SE discussions but often fail to account for contextual nuances. Early SE toxicity detection often relied on lexicon-based approaches or classical machine learning algorithms like SVMs. Lexicon-based approaches use predefined dictionaries of words or phrases associated with specific sentiments or toxicity levels. For example, terms like ”kill” or ”junk” might be flagged as toxic based on such a dictionary, without considering their contextual meaning [28]. These methods, while easy to implement and understand, often misinterpret technical jargon or colloquial expressions common in SE communication, leading to a high rate of false positives [46]Raman et al. [41] demonstrated the use of an SVM tailored for GitHub issue comments, revealing its potential for identifying toxicity trends across communities, but its performance was limited by its reliance on simple feature sets. Moreover, empirical evaluations have shown that while lexicon-based models can achieve acceptable precision on datasets like Code Review and Gitter Ethereum, their recall remains poor when handling nuanced or domain-specific expressions [44, 46].
@@
In contrast, LLM-based models capitalize on generative pre-training to develop richer contextual understanding. These methods utilize transformer-based architectures (BERT, RoBERTa, GPT-3, GPT-4), trained on large-scale, generative pre-training objectives. For example, generative pre-training involves training models to predict masked words or the next token in a sequence, enabling them to capture rich semantic relationships in text [30, 51]. ToxiSpanSE [47], a fine-tuned RoBERTa-based model, achieved an impressive F1-score of 0.88 in detecting toxic spans in SE datasets, significantly outperforming earlier lexicon-based methods. Similarly, Shafikuzzaman et al. [48] evaluated GPT-3.5 and GPT-4, demonstrating their strong performance in zero-shot and few-shot settings on SE datasets like ARC and GIT. Another innovative example is the DeMod tool [31], which incorporates explainable AI capabilities to enhance both detection and mitigation of toxicity. By leveraging ChatGPT’s generative pre-training, DeMod offers fine-grained explanations and personalized modification suggestions to address toxic content effectively.
@@ 3.3.3 Task Type
Most toxicity detection approaches focus on a binary classification task, labeling each comment as either toxic or non‑toxic. For instance, the five datesets from 4 are all examples of binary-classified datasets from the SE community. Most models built using these datasets yield an overall yes/no decision without identifying which parts of the text are problematic. Models built using these datasets typically yield an overall yes/no decision without pinpointing which parts of the text are problematic. In contrast, one notable model adopts span‑level detection methods that identify the exact segments containing toxic language, thereby increasing interpretability and transparency [47].
Recent work has also explored multiclass classification to capture different levels of toxicity within a single input, offering an alternative to the conventional binary approach. For instance, one study focusing on maternal insults and stereotyping classified texts into multiple categories—such as “contains maternal insult,” “contains slight maternal insult,” “hard to say,” or “does not contain maternal insult”—and reported higher accuracy compared to binary classification [50]. However, the decision boundaries of these multiclass models are not entirely transparent. Moreover, settings such as the model’s temperature can also impact the results, leading to variability even when using the same prompt.
Beyond detection, several studies have investigated generative mitigation strategies. For example, a Conflict Reduction System (CRS) has been proposed that not only detects offensive language but also generates rephrased alternatives to reduce hostility in communities such as GitHub, Gitter, Slack, and Stack Overflow [13]. Similarly, the DeMod tool leverages explainable AI to provide fine‑grained detection results along with personalized content modification suggestions, thereby helping users mitigate toxicity before posting [31]. Despite its innovative nature, this approach remains in its early stages and requires further empirical validation on diverse and representative datasets.
@@ 4 Future Directions
As research on toxicity detection evolves, several key areas emerge as promising directions for future exploration. Advancements in preprocessing techniques can enhance data quality while preserving critical contextual nuances, improving model performance. Similarly, refining detection methods by integrating external knowledge and conversational context may help address challenges in accurately interpreting toxicity. Additionally, improving model interpretability and transparency through explainable AI techniques can foster trust and usability in automated toxicity detection systems. This section discusses these potential advancements, outlining future opportunities in preprocessing, detection, and explainability to guide ongoing research in this domain.
4.1 Preprocessing
@@
A potential avenue for future work lies in refining data cleaning techniques to better preserve nuanced semantic meanings in text. For instance, while standard normalization processes often reduce elongated words such as ”loooove” to their base form ”love,” this approach may overlook subtle variations in emotional intensity or intent, potentially leading to the loss of valuable contextual or toxicity-related information. Similarly, symbol removal practices might benefit from greater specificity; instead of blanket removal, future research could explore methods to distinguish between symbols that contribute to textual semantics and those that introduce noise. Developing more granular preprocessing strategies that account for these subtleties could enhance the ability of toxicity detection models to capture contextually rich and emotionally complex patterns, ultimately improving their accuracy and robustness.
4.2 Detection
@@
A major challenge in toxicity detection is accurately interpreting context, as certain phrases may be toxic in one setting but neutral or even positive in another. Future work could explore incorporating external knowledge sources, conversation history, or user intent modeling to enhance contextual understanding. For example, leveraging transformer-based models with multi-turn dialogue context could help distinguish between sarcasm, humor, and genuinely harmful content. A recent research [4] demonstrates that human perception of toxicity is highly dependent on preceding dialogue, and integrating this context into detection models significantly improves accuracy, highlighting the importance of conversational context in toxicity detection. Zhu et al. [61] present a framework that integrates topic modeling with external knowledge bases, offering a potential direction for improving toxicity classification by leveraging similar techniques to understand user intent and conversational context. Future research could build upon these findings by combining multi-turn dialogue modeling, knowledge-aware transformers, and socio-linguistic analysis to develop more nuanced toxicity detection systems capable of differentiating between targeted toxicity and benign discussions.
@@
Many toxicity detection models function as black boxes, making it challenging for users and moderators to understand why content is flagged. To improve transparency, future research should prioritize explainable AI (xAI) techniques, such as attention visualization and SHAP (SHapley Additive exPlanations) [32]. SHAP assigns contribution scores to individual features, helping to identify which words or phrases influence a model’s toxicity predictions. By quantifying these contributions, it enhances interpretability and allow researchers to assess whether model decisions align with human judgment. Recent models increasingly emphasize interpretability; for example, ToxiSpanSE [47] improves transparency by detecting toxic spans in code review comments and highlighting specific toxic segments. Additionally, human-in-the-loop approaches that enable users to contest or refine model decisions could foster trust and usability while mitigating unintended consequences, such as excessive or insufficient content flagging.
@@ 5 Ablation Study
To evaluate the specific contributions of different prompting strategies and model choices, we conducted an ablation study. The goal was to measure their individual impacts on toxicity detection and mitigation tasks within SE discussions. The following sections describe our experimental setups and findings.
5.1 Experiments
Two experiments will be presented below: one for toxicity detection and another for toxicity mitigation. The detection experiment assessed the performance of LLMs in classifying toxic GitHub issues, while the mitigation experiment focused on their capability to rewrite individual toxic comments.
5.1.1 Experiment Setup for Toxicity Detection
@@
To empirically evaluate the feasibility of leveraging LLMs for toxicity detection in SE communities, we conducted experiments using a dataset of 194 GitHub issues. Each issue was annotated as either containing toxic content or not. This dataset was created and shared by Ramen. In the original study, the authors trained an SVM classifier and tested their model on the same dataset.
The GitHub issues were collected using the GitHub API and GHTorrent and were manually labeled by the original authors. We evaluated two LLMs — GPT-4o and Gemini-1.5-flash — under two prompting conditions. In the few-shot setting, we selected three randomly chosen toxic examples and three non-toxic examples from the original paper’s training set as prompts. In the zero-shot setting, only a textual definition of toxicity was provided without any examples. After obtaining each model’s predictions (“Yes” for Toxic and “No” for Non-Toxic), we computed classification metrics, including accuracy, precision, recall, and F1-score for both classes, as well as overall accuracy.
5.1.2 Experiment for Toxicity Mitigation
@@
In the mitigation experiment, we continued utilizing the dataset from the same source. The key difference from the detection experiment is that, instead of entire issues, the input consisted of individual comments. This granularity enables a more precise evaluation of the mitigation effects [41]. As illustrated in Figure 5, we employed a structured approach to ensure model independence by utilizing three distinct models, preventing over-reliance on any single approach and enabling a more robust mitigation strategy. Our pipeline first invokes a few-shot GPT-4o, selected due to its exceptional performance in the previous detection experiment 5.1.1, to identify if a comment is toxic. Prior studies have shown that large-scale language models such as GPT-4 outperform traditional classifiers in zero-shot and few-shot settings, particularly in identifying offensive and harmful language while maintaining high recall rates [48]. The ChatGPT series of models has been widely recognized for its effectiveness in handling nuanced toxicity detection, especially in software engineering discussions [34]. Then, the comment will be sent to the Gemini-1.5-flash model to rewrite the text as it has demonstrated strong performance in text generation while preserving semantic consistency. Unlike rule-based text sanitization methods, generative models like Gemini-1.5-flash provide more contextually appropriate rephrasings, ensuring that the rewritten text remains meaningful within the software engineering domain while effectively reducing toxicity [50]. Finally, to evaluate the effectiveness of the rewrites, we measured the semantic similarity using Sentence-BERT, a model designed for capturing sentence-level meaning beyond surface-level wording [48]. Sentence-BERT was selected due to its superior ability to assess the semantic consistency of rewritten comments compared to classical cosine similarity approaches, making it particularly effective for our mitigation pipeline [41]. This approach provided a robust estimation of how well the mitigated comments preserved the intent of the initial submissions rather than merely matching lexical choices.
Figure 5: Mitigation Experiment Workflow
@@
5.2 Results

### v2

@@ 3.3.2 Architectural difference
Traditional models offer a straightforward approach to toxicity detection in SE discussions but often fail to account for contextual nuances. Early SE toxicity detection often relied on lexicon-based approaches or classical machine learning algorithms like SVMs. Lexicon-based approaches use predefined dictionaries of words or phrases associated with specific sentiments or toxicity levels. For example, terms like “kill” or “junk” might be flagged as toxic based on such a dictionary, without considering their contextual meaning [jongeling2017negative]. These methods, while easy to implement and understand, often misinterpret technical jargon or colloquial expressions common in SE communication, leading to a high rate of false positives [sarker2020benchmark]raman2020stress demonstrated the use of an SVM tailored for GitHub issue comments, revealing its potential for identifying toxicity trends across communities, but its performance was limited by its reliance on simple feature sets. Moreover, empirical evaluations have shown that while lexicon-based models can achieve acceptable precision on datasets like Code Review and Gitter Ethereum, their recall remains poor when handling nuanced or domain-specific expressions [sarker2022identification, sarker2020benchmark].
In addition to traditional models like lexicon, some attention-based methods have also been applied to the field of toxicity detection. These neural models explicitly learn to weigh tokens or sentences by importance, typically adding word- or sentence-level attention atop lightweight encoders (e.g., GRU or CNN–RNN hybrids) to spotlight toxic cues. pavlopoulos2017deeper showed that RNN based on word vectors outperforms logistic regression or MLP classifiers with character or word n-grams in review. At the same time, this method also brings another benefit: suspicious words can be highlighted for free without including labeled words in the training data. Subsequent studies compared contextual vs. self-attention and confirmed the benefit of contextual attention for implicit abuse [chakrabarty2019pay]. Recently, an attention-enhanced CNN–LSTM achieved an accuracy of 0.76 on the shared task on Abusive Comment Detection in Tamil - ACL 2022, outperforming deep learning models like CNN, LSTM, Bi-LSTM  [first2024attention].
@@
LLM-based models capitalize on generative pre-training to develop richer contextual understanding. These methods utilize transformer-based architectures (BERT, RoBERTa, GPT-3, GPT-4), trained on large-scale, generative pre-training objectives. For example, generative pre-training involves training models to predict masked words or the next token in a sequence, enabling them to capture rich semantic relationships in text [kenton2019bert, tan2022roberta]. ToxiSpanSE [sarker2023toxispanse], a fine-tuned RoBERTa-based model, achieved an impressive F1-score of 0.88 in detecting toxic spans in SE datasets, significantly outperforming earlier lexicon-based methods. Similarly, shafikuzzaman2024empirical evaluated GPT-3.5 and GPT-4, demonstrating their strong performance in zero-shot and few-shot settings on SE datasets like ARC and GIT. Another innovative example is the DeMod tool [li2024demod], which incorporates explainable AI capabilities to enhance both detection and mitigation of toxicity. By leveraging ChatGPT’s generative pre-training, DeMod offers fine-grained explanations and personalized modification suggestions to address toxic content.
In order to compare the commonly used transformer-based models and their phenotypes for toxicity detection, we conducted an experiment. Five Transformer variants are fine tuned under a uniform recipe (1 epoch, batch size 32, learning rate 2e-5, max sequence length 512, using an 80/20 split for the training and testing datasets) [devlin2019bert, liu2019roberta, trautsch2022predicting]. The dataset on software engineering code reviews is shared by [sarker2023automated]. Table 10 summarizes each model’s architecture and classification performance. Among the base‐size Transformers, seBERT—built on the BERT-base backbone with SE-specific tokenization—achieves the highest accuracy (0.9463) and F1 score (0.9461). BERT-base follows closely (0.9456 accuracy, 0.9450 F1), while RoBERTa-base trails slightly (0.9420, 0.9414). Scaling up to the large variants under a single‐epoch, default‐hyperparameter regime yields no clear gains: BERT-large simply matches BERT-base (0.9456, 0.9450), and RoBERTa-large actually performs worse (0.9395, 0.9385). These results indicate that for mid-sized SE toxicity datasets, base-size models strike a better balance of capacity and stability. Larger models—with doubled depth and hidden dimensions—tend to underfit or train unstably when constrained to one epoch; fully leveraging their extra capacity therefore requires stronger optimization momentum, more training data, and additional epochs to achieve further performance improvements.
Table 10: Architecture and performance of BERT-family models
Model Parameters Layers Hidden Size Heads Embedding Size Accuracy F1
BERT-base 110 M 12 768 12 768 0.9456 0.9450
BERT-large 340 M 24 1024 16 1024 0.9456 0.9450
RoBERTa-base 125 M 12 768 12 768 0.9420 0.9414
RoBERTa-large 355 M 24 1024 16 1024 0.9395 0.9385
seBERT 110 M 12 768 12 768 0.9463 0.9461
Despite their impressive F1 scores, transformer-based toxicity detectors inherit two well-documented weaknesses. First, they are highly sensitive to adversarial perturbations that confuse malicious tokens through character-level perturbations. The detection recall rate of BERT’s detector will be reduced by more than 50% in some cases [kurita2019towards], and recent benchmarks such as TaeBench record attack-success rates above 77% against both open-source and commercial APIs [zhu2024taebench]nozza2023state also showed that the various “fancy writing” of swear words/abusive words is enough to significantly reduce the recall rate of mainstream transformer-based detectors. Second, detectors’ performance degrades sharply under domain shift. Empirical studies show that transformer models fine-tuned on the Jigsaw/Wikipedia Toxic Comment corpus lose 40–53% of their F1 when applied to software-engineering data such as GitHub code review or Gitter messages [sarker2020benchmark]. The other complementary cross-platform analysis finds that fine-tuned language models keep only around 50% of their in-domain performance under cross-domain testing [schouten2023cross].
In response to these shortcomings, an open source tool PROF was developed, which is used to automatically generate a variety of “fancy writing methods” and inject them into the training set, significantly improving the robustness of the model under such disturbances in a confrontational data enhancement manner [nozza2023state].
@@ 3.3.3 Task type
Most toxicity detection approaches focus on a binary classification task, labeling each comment as toxic or non‑toxic. For instance, the five datasets from Table 6 are all examples of binary-classified datasets from the SE community. Models built using these datasets typically yield an overall yes/no decision without pinpointing which parts of the text are problematic.
In contrast, span-level detection identify the exact tokens that carry toxic meaning. Span-based tools such as ToxiSpanSE highlight the problematic words inside code review comments, giving moderators an immediate rationale for the model’s decision [sarker2023toxispanse]. Table 11 summarizes the detection performance of ToxiCR and ToxiSpanSE. The data is selected from their fine-tuned model based on BERT, and they use the same Code Review dataset. From the data, ToxiCR slightly outperforms in overall classification. However, ToxiSpanSE provides the added benefit of fine-grained localization of toxic tokens. In practice, the two can be combined: first apply ToxiCR for high-throughput filtering, then invoke ToxiSpanSE on flagged comments to obtain precise, explainable toxic spans—thereby balancing efficiency with interpretability.
Table 11: Performance comparison between ToxiCR and ToxiSpanSE. Data is from sarker2023automated, sarker2023toxispanse.
Tool Task Class Precision Recall F1-Score
ToxiCR Comment-level binary Toxic 0.907 0.874 0.889
Non-Toxic 0.970 0.978 0.974
ToxiSpanSE Token-level span det. Toxic words 0.870 0.890 0.860
Non-Toxic words 0.950 0.950 0.950
Recent work has also explored multiclass classification to capture different levels of toxicity within a single input, offering an alternative to the conventional binary approach. For instance, one study focusing on maternal insults and stereotyping classified texts into multiple categories—such as “contains maternal insult”, “contains slight maternal insult”, “hard to say”, or “does not contain maternal insult”—and reported higher accuracy compared to binary classification [sultana2024exploring]. However, the decision boundaries of these multiclass models are not entirely transparent. Moreover, settings such as the model’s temperature can also impact the results, leading to variability even when using the same prompt.
The choice of task granularity directly affects human review effort. Whereas binary tags force reviewers to reread the entire article to find the problem, span tags allow them to jump directly to the offending passage. Several user studies with reviewers have shown that better interpretability leads to improved efficiency and accuracy [xiang2021toxccin, song2023modsandbox]. However, currently there are no controlled user studies in software engineering to quantify how span-level, sentence-level, or binary outputs affect reviewer efficiency and decision quality at scale. Therefore, future work could include user comparison studies to measure average processing time per review in different output formats, decision accuracy with and without model support, cognitive burden, and trust in the system.
Beyond detection, several studies have investigated generative mitigation strategies. For example, a Conflict Reduction System (CRS) has been proposed that not only detects offensive language but also generates rephrased alternatives to reduce hostility in communities such as GitHub, Gitter, Slack, and Stack Overflow [cheriyan2021towards]. Similarly, the DeMod tool leverages explainable AI to provide fine‑grained detection results along with personalized content modification suggestions, thereby helping users mitigate toxicity before posting [li2024demod]. Despite its innovative nature, this approach remains in its early stages and requires further empirical validation on diverse and representative datasets.
@@ 4 Future directions
Future research in toxicity detection could focus on several key areas: enhancing pre-processing techniques, refining detection methods with contextual understanding, improving model interpretability and explainability, and critically addressing ethical governance and policy implications. The following sections outline opportunities within these domains to guide ongoing research.
4.1 Context-preserving pre-processing
@@
A potential avenue for future work lies in refining data cleaning techniques to better preserve nuanced semantic meanings in text. For instance, while standard normalization processes often reduce elongated words such as “loooove” to their base form “love”, this approach may overlook subtle variations in emotional intensity or intent, potentially leading to the loss of valuable contextual or toxicity-related information. Similarly, symbol removal practices might benefit from greater specificity; instead of blanket removal, future research could explore methods to distinguish between symbols that contribute to textual semantics and those that introduce noise. This could involve developing symbol-aware tokenizers that apply language-specific heuristics, such as regex-based detection of emoticons, prior to general symbol removal. Such an approach enables the nuanced preservation of sentiment-conveying symbols, which can then be effectively integrated into analysis, as demonstrated by methods like the blending approach highlighted by [zou2022sentiment] for enhancing semantic understanding. Developing more granular pre-processing strategies that account for these subtleties could enhance the ability of toxicity detection models to capture contextually rich and emotionally complex patterns, ultimately improving their accuracy and robustness.
4.2 Context-aware detection methods
@@
A major challenge in toxicity detection is accurately interpreting context, as certain phrases may be toxic in one setting but neutral or even positive in another. Future work could explore incorporating external knowledge sources, conversation history, or user intent modeling to enhance contextual understanding. For example, leveraging transformer-based models with multi-turn dialogue context could help distinguish between sarcasm, humor, and genuinely harmful content. A recent research [anuchitanukul2022revisiting] demonstrates that human perception of toxicity is highly dependent on preceding dialogue, and integrating this context into detection models significantly improves accuracy, highlighting the importance of conversational context in toxicity detection. zhu2021topic present a framework that integrates topic modeling with external knowledge bases, offering a potential direction for improving toxicity classification by leveraging similar techniques to understand user intent and conversational context. Future research could build upon these findings by combining multi-turn dialogue modeling, knowledge-aware transformers, and socio-linguistic analysis to develop more nuanced toxicity detection systems capable of differentiating between targeted toxicity and benign discussions.
4.3 Model interpretability and explainability
@@
Many toxicity detection models function as black boxes, making it challenging for users and moderators to understand why content is flagged. To improve transparency, future research should prioritize explainable AI (xAI) techniques, such as attention visualization and SHAP (SHapley Additive exPlanations) [lundberg2017unified]. SHAP assigns contribution scores to individual features, helping to identify which words or phrases influence a model’s toxicity predictions. By quantifying these contributions, it enhances interpretability and allow researchers to assess whether model decisions align with human judgment. Recent models increasingly emphasize interpretability; for example, ToxiSpanSE [sarker2023toxispanse] improves transparency by detecting toxic spans in code review comments and highlighting specific toxic segments. Additionally, human-in-the-loop approaches that enable users to contest or refine model decisions could foster trust and usability while mitigating unintended consequences, such as excessive or insufficient content flagging.
4.4 Ethical and Policy Considerations
Beyond technical advancements, future directions must critically address the ethical governance [olsen2024right] and policy implications of deploying toxicity detection systems in real-world SE platforms. This includes investigating the development of clear community moderation guidelines that align with platform values and user expectations, ensuring transparency in automated decision-making [tiwari2023explainable]. Research should also explore robust user feedback mechanisms, allowing individuals to contest classifications and contribute to the refinement of detection models [molina2022ai]. Furthermore, it is crucial to analyze the broader policy frameworks surrounding content moderation, user consent, and data privacy to ensure that technical enhancements are implemented responsibly and uphold user rights and ethical standards [llanso2020artificial]. This holistic approach is essential for fostering trust and ensuring the responsible application of toxicity detection technologies.
@@ 5 Ablation study
To measure the individual impact of prompting strategies and model choices on SE discussions, we conducted an ablation study with two experiments. The first focused on toxicity detection (classifying code review comments), and the second focused on toxicity mitigation (rewriting individual toxic comments).
5.1 Detection experiment
@@
To empirically evaluate the feasibility of leveraging LLMs for toxicity detection in SE communities, we conducted experiments using a dataset of 19,651 labeled code review comments. Each issue was annotated as either toxic or non-toxic. About 19% of the comments are labeled as toxic. This dataset was created and shared by  sarker2023automated. In the original study, the authors developed a supervised learning based tool to identify toxic code review comments.
The code review comments were collected from FOSS project code reviews and Gitter chat through a four-step process: data mining, stratified sampling, manual labeling, dataset aggregation. We chose three LLMs: GPT-4.1, Gemini-1.5-flash and deepseek-chat. We selected these models due to robust API support and documentation from their providers. Both Google, OpenAI and DeepSeek provide convenient API interfaces and detailed documentation. Among the OpenAI models, GPT-4.1 understands complex instructions and its I/O cost (2/2/8 per million tokens) is 26% lower than GPT-4o [openai2024gpt41]. The deepseek-chat model is selected due to its open-source nature and ability to be deployed on local infrastructure, a feature that is particularly valuable for organizations concerned with data privacy or operational costs. Our goal is to assess whether it can deliver toxicity detection performance on par with more resource-intensive proprietary models.All LLMs are evaluated under two prompting conditions. In the few-shot setting, we selected three randomly chosen toxic examples and three non-toxic examples from a different dataset built by raman2020stress to maintain independence. In the zero-shot setting, only a textual definition of toxicity was provided without any examples. After obtaining each model’s predictions (“Yes” for Toxic and “No” for Non-Toxic), we computed classification metrics, including accuracy, precision, recall, and F1-score for both classes, as well as overall accuracy.
5.2 Mitigation experiment
@@
In the mitigation experiment, we continued utilizing the dataset from the same source. The key difference from the detection experiment is that, instead of entire issues, the input consisted of individual comments. This granularity enables a more precise evaluation of the mitigation effects [raman2020stress]. As illustrated in Figure 5, we employed a structured approach to ensure model independence by utilizing four distinct models, preventing over-reliance on any single approach and enabling a more robust mitigation strategy. Our pipeline first invokes a few-shot GPT-4.1, selected due to its exceptional performance in the previous detection experiment, to identify if a comment is toxic. Prior studies have shown that large-scale language models such as GPT-4 outperform traditional classifiers in zero-shot and few-shot settings, particularly in identifying offensive and harmful language while maintaining high recall rates [shafikuzzaman2024empirical]. The ChatGPT series of models has been widely recognized for its effectiveness in handling nuanced toxicity detection, especially in software engineering discussions [mishra2024exploring].
Then, the comment will be sent to the Gemini-1.5-flash model and deepseek-chat model to rewrite the text as it has demonstrated strong performance in text generation while preserving semantic consistency. Unlike rule-based text mitigation methods, generative models like Gemini-1.5-flash and deepseek-chat model provide more contextually appropriate rephrasing, ensuring that the rewritten text remains meaningful within the software engineering domain while effectively reducing toxicity [sultana2024exploring].
Finally, to evaluate the effectiveness of the rewrites, we measured the semantic similarity using Sentence-BERT, a model designed for capturing sentence-level meaning beyond surface-level wording [shafikuzzaman2024empirical]. Sentence-BERT was selected due to its superior ability to assess the semantic consistency of rewritten comments compared to classical cosine similarity approaches, making it particularly effective for our mitigation pipeline [raman2020stress]. This approach provided a robust estimation of how well the mitigated comments preserved the intent of the initial submissions rather than merely matching lexical choices.
Figure 5: Experimental workflow for toxicity detection and mitigation.
@@
5.3 Prompt and hyper-parameter specification
To remove any ambiguity about reproduction, we used the following prompt wording, model releases, and decoding settings used in every experiment.
Zero-shot prompt: You are a toxicity detection assistant. In this task, you will read a code review comment. If any part contains rude, offensive, or disrespectful language, respond with ’Yes’. Otherwise, ’No’. Comment: <comment>
The few-shot prompt used the same introductory lines, then appended six fixed exemplars—three toxic and three non-toxic—sampled with random state of 42. The exemplar block, is stored in a CSV file once generated, is identical for every run, after which the line Comment: comment inserts the test instance.
All calls were issued via the official APIs: OpenAI GPT-4.1, Google Gemini-1.5-flash and DeepSeek deepseek-chat. We did not use stop sequences, presence or frequency penalties, or logit bias, and each request returned a single completion (n = 1). Table 12 lists the full decoding settings.
Table 12: Hyper-parameter specification of LLM variants evaluated in experiments.
Task Model Temperature Top-p Top-k Max tokens Other
Detection GPT-4.1 1.0 1.0 2 048
Detection Gemini-1.5-flash 1.0 1.0 40 2 048
Detection deepseek-chat 1.0 1.0 50 4 096 128K context length
Mitigation Gemini-1.5-flash 1.0 1.0 40 2 048
Mitigation deepseek-chat 1.0 1.0 50 4 096 128K context length
Semantic similarity SBERT (all-mpnet-base-v2) n/a n/a n/a 512 768-d embeddings
6 Results